2m+n=12
m-9n=25
find the substitution method

Answer by rfer(16322)

m=9n+25
2m+n=12
2(9n+25)+n=12
18n+50+n=12
19n=12-50
19n=-38
n=-2
--------------------
2m+-2=12
2m=14
m=7
---------------
7-9*-2=25
7+18=25